[figure: 9.7 cm]

MAGNETIC TRANSITION IN La_{1/3}Nd_{1/3}Ca_{1/3}MnO_{3} PEROVSKITE INDUCED BY ELECTRIC FIELD AT A CONSTANT MAGNETIC FIELD

J. Baszyński^{a}, A. Kowalczyk^{a} and J. Kováč^{b}

aInstitute of Molecular Physics, Polish Academy of Sciences, Smoluchowskiego 17,

60-179 Poznań, Poland

bInstitute of Experimental Physics, SAS, Watsonova 47, 043-53 Košice, Slovakia

We report a discovery of a new effect in La_{1/3}Nd_{1/3}Ca_{1/3}MnO_{3} – the jump of magnetization induced by external electric field E (2.5 – 4) V/cm at a given magnetic field H (18 – 25) kOe at T = 4.2 K. This effect is large when compared with a metamagnetic transition induced by the magnetic field in many substances. The origin of the magnetization jump induced by the external electric field has not yet been explained. We assume that this transition is induced by the hopping of E-excited e_{g} electrons and that this is a source of a jumpwise increase in magnetization in the noncollinear magnetic phase within the perovskite. This effect can be attributed to a strong coupling between carriers induced by the electric field (the conduction electrons e_{g}), and the local magnetic moments (t_{2g} – localized spins) through a mechanism the Zener double exchange.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : J. Kováč

Address for correspondence : Institute of Experimental Physics, Watsonova 47, 043-53 Košice, Slovakia

Email address : jkovac@kosice.upjs.sk